# Rethinking MUSHRA: Addressing Modern Challenges in Text-to-Speech Evaluation

### v1

@@ Abstract
Despite rapid advancements in TTS models, a consistent and robust human evaluation framework is still lacking. For example, MOS tests fail to differentiate between similar models, and CMOS’s pairwise comparisons are time-intensive. The MUSHRA test is a promising alternative for evaluating multiple TTS systems simultaneously, but in this work we show that its reliance on matching human reference speech unduly penalises the scores of modern TTS systems that can exceed human speech quality. More specifically, we conduct a comprehensive assessment of the MUSHRA test, focusing on its sensitivity to factors such as rater variability, listener fatigue, and reference bias. Based on our extensive evaluation involving 471 human listeners across Hindi and Tamil we identify two primary shortcomings: (i) reference-matching bias, where raters are unduly influenced by the human reference, and (ii) judgement ambiguity, arising from a lack of clear fine-grained guidelines. To address these issues, we propose two refined variants of the MUSHRA test. The first variant enables fairer ratings for synthesized samples that surpass human reference quality. The second variant reduces ambiguity, as indicated by the relatively lower variance across raters. By combining these approaches, we achieve both more reliable and more fine-grained assessments. We also release Mango, a massive dataset of 47,100 human ratings, the first-of-its-kind collection for Indian languages, aiding in analyzing human preferences and developing automatic metrics for evaluating TTS systems.
@@ 1 Introduction
The above problem is well studied for the Mean Opinion Scores (MOS) test (Wester et al., 2015; Finkelstein et al., 2023; Kirkland et al., 2023) which has received much constructive criticism over the past few years. Specifically, in a MOS test, listeners assess each system independently, which can result in an inability to accurately capture the subtle relative differences between similar systems. This poses a significant challenge in modern TTS evaluation where systems that perform equally well need to be compared against each other. To address these issues some of the recent works rely on CMOS tests (Loizou, 2011). However, this test is costly and time-consuming as it involves ($N\choose 2$ comparisons between all pairs of $N$ systems.
The MUSHRA test has gained popularity in addressing these issues. This test scales better by enabling a parallel comparison of the $N$ systems, and addresses the limitations of MOS tests that only allow isolated evaluation. However, we show that even the MUSHRA test is not devoid of issues. To begin with, we note that the MUSHRA test was conventionally designed to assess intermediate-quality audio systems (ITU-R, 2015). However, state-of-the-art TTS systems (Ju et al., 2024) are not of intermediate quality and instead generate audios having quality on par or even better than human recordings. To align with these modern developments, several works adopt variants of MUSHRA (Merritt et al., 2022; Li et al., 2023; Shen et al., 2024), which differ in implementation but the validity of these modified tests is unknown.
@@
With the goal of seeking answers to the above questions, we collected 47,100 human ratings by conducting the MUSHRA test involving 3 systems across two languages, viz., Tamil and Hindi. Our in-depth analysis based on these ratings, reveals two primary shortcomings: (i) reference-matching bias and (ii) judgement ambiguity. To mitigate these issues, we propose two refined variants of the MUSHRA test. The first variant does not explicitly identify the human reference to the rater. Doing so, prevents unfair penalties for well-synthesized samples that differ from the human reference, such as those with natural prosody that do not match the reference’s prosody. In the second variant, raters are provided scoresheets to systematically calculate MUSHRA scores, by explicitly marking pronunciation mistakes, unnatural pauses, digital artifacts, word skips, liveliness, voice quality, rhythm, etc. Using the scores for these fine-grained criteria, they arrive at the final MUSHRA score. Our studies show that both these variants lead to a more reliable evaluation with the second variant also allowing for fine-grained fault isolation during evaluation. We then show that a combination of these two approaches that leverages their individual strengths ensures both consistency and granularity. It allows modern TTS systems to be evaluated without being unfairly penalized for surpassing the reference in naturalness or prosody. The detailed scoring for pronunciation, prosody, and other factors provides actionable insights,and helps practitioners understand precisely where their systems excel and where improvements are needed. This combination creates a more balanced and sensitive evaluation framework, offering a clearer and more reliable assessment of TTS system performance.
@@
    Large-scale empirical validation of proposed variants resulting in Mango, a dataset of 47,100 ratings from 471 listeners across Hindi and Tamil, examining three TTS systems.
@@ 3 MANGO: A Corpus of Human Ratings for Speech
We introduce a new dataset, Mango: MUSHRA Assessment corpus using Native listeners and Guidelines to understand human Opinions at scale. It is a first-of-its-kind collection for any Indian language, comprising 47,100 human ratings of TTS systems and ground-truth human speech in both Hindi and Tamil. Given the shortcomings of Mean Opinion Score (MOS) and Comparative Mean Opinion Score (CMOS) tests, our goal is to critically examine a promising alternative—the MUSHRA test—by conducting a large-scale evaluation involving multiple raters, systems, and languages. To do so, we adopt the standard MUSHRA test (ITU-R, 2015). Raters evaluate multiple stimuli on each page, including an explicit (mentioned) that serves as a benchmark for high-quality speech, along with an anchor and implicit (hidden) reference to calibrate judgments. Each stimulus is rated on a continuous scale from 0 to 100, which is also discretized: 100-80 (Excellent), 80-60 (Good), 60-40 (Fair), 40-20 (Poor), and 20-0 (Bad). We describe our evaluation setup below, and provide the detailed instructions provided to participants in Appendix A.1.
@@ 3.2 Synthesizing Speech Samples for Annotation
To generate samples for TTS evaluation, we train TTS systems on the Hindi and Tamil subsets of the IndicTTS database (Baby et al., 2016). Each language consists of recordings from a female and male speaker (Hindi: 20.17 hours; Tamil: 20.59 hours). We train FastSpeech2 (FS2) (Ren et al., 2021) with HiFiGAN v1 (Kong et al., 2020), VITS (Kim et al., 2021) from scratch on the train-test splits using hyper-parameters suggested in a recent study (Kumar et al., 2023). We finetune StyleTTS2 (ST2) (Li et al., 2023) from the LibriTTS checkpoint and XTTSv2 (CoquiAI, 2023) from the multilingual checkpoint with the hyperparameters from their original implementations on the same splits.
@@ 3.3 Annotation Process and Dataset Statistics
With the above process we collected 47,100 human ratings for TTS systems. Table 1 shows the demographic distribution, the number of participants and the overall number of ratings across all MUSHRA tests and our proposed variants (MUSHRA-NMR, MUSHRA-DG, MUSHRA-DG-NMR) which are described in Section 5.
@@ 4.1 Is MUSHRA A Reliable Test?
In Table 3, we present the results of the MUSHRA test among 3 systems and find VITS and ST2 score highest in Hindi and Tamil respectively. Surprisingly, all systems attain scores in the “Good" bin with MUSHRA scores between 60 and 80, while the reference surpasses all systems with scores in the “Excellent" bin. Given that state-of-the-art TTS systems are able to reach quality on par with references, one would expect a much smaller gap between the reference and systems. To confirm this, we conduct the more reliable but expensive CMOS test with 15 listeners in each language. In this test, we ask the rater to compare a given system, such as VITS, with a reference audio sample. The rater evaluates both the reference and the output from a system being tested without prior knowledge of which audio sample corresponds to which system, ensuring an unbiased comparison. The raters assign a single score ranging from -3 to +3 in increments of 0.5. A score of -3 indicates that System A is much worse than System B . A score of +3 indicates that System A is much better than System B. A score of 0 means that both systems are equal in quality. As seen from the scores in Table 3, CMOS indicates that the outputs synthesised by VITS and ST2 are very close in quality to the reference in Hindi and Tamil respectively, while MUSHRA scores do not reflect this at all. We hypothesize that listeners in the MUSHRA test are subject to various biases, one of which we term the reference-matching bias. This bias may lead to situations where systems that perform comparably to or better than the reference are rated less favorably, as listeners tend to focus on aligning their ratings with the reference outputs while evaluating the systems. While this may have been acceptable when TTS systems lagged behind human speech quality, it is undesirable in the current scenario where modern TTS systems often exceed the reference in aspects like naturalness and prosody. This suggests that the MUSHRA test, in its conventional form, may no longer be sufficient for evaluating state-of-the-art TTS systems. Instead, alternative methodologies, such as the variants we propose in Section 5, may help ensure more fair and accurate assessments.
@@ 4.2 How reliable is the mean statistic in MUSHRA scores?
To delve deeper into judgment ambiguity, we examine variations between two systems. We consider an utterance where the mean scores for the samples generated by VITS and ST2 are nearly identical, but the variance across raters for each system is high. This high variance indicates significant ambiguity. Upon listening to many such utterances and speaking to many raters, we hypothesize that the ambiguity likely stems from different raters focusing on different aspects of the generated samples. For instance, some raters may prioritize prosody, others voice quality, and yet others the presence of digital artifacts. We hypothesize that asking raters to highlight these subtle differences across multiple dimensions while assigning a single score can lead to ambiguity in determining how much to penalize or reward a system’s score. Hence, clear guidelines which take into account a fine-grained evaluation across different aspects would help (as proposed later in Section 5).
@@ 4.4 What is the impact of rejecting raters per standard MUSHRA protocol?
Traditionally, MUSHRA employs a rater rejection criterion, wherein raters scoring the hidden reference (HR) below a threshold ($\lambda$) more than 15% of the time are rejected. This rejection rule stems from the inherent assumption that the HR is the gold standard, which is not true in modern TTS settings where TTS systems (Li et al., 2023) claim to achieve performance surpassing the reference. In such cases, a rater consistently scoring the HR lower might not necessarily reflect unreliability, but rather a nuanced perception of the reference’s limitations compared to the evaluated systems. This is reinforced by observations from Table 3 where raters clearly prefer ST2 over reference for Tamil with a mean CMOS score of 0.24. This observation is further supported by the MUSHRA scores shown in Figure 3. The table shows that while rejecting raters based on the conventional MUSHRA criterion does not affect system rankings, it does notably shift scores. Specifically, system scores decrease with increasing $\lambda$, while reference scores increase. This trend hints at the reference-matching bias wherein raters who give the HR high scores might be unconsciously matching system samples to the mentioned reference, rather than rating based on their absolute perception of quality.
@@ 4.6 Does adding more systems affect scores?
To better understand cognitive overload in the MUSHRA test, we scale up the number of systems to be rated by introducing one new competitive system - XTTS, and repeating an existing system - VITS (VITS-R) in the original Hindi MUSHRA test. We call this MUSHRA-Extended. The results show that raters were highly consistent, with VITS and VITS-R receiving nearly identical scores (68.99 and 68.47, respectively), despite the randomized order. Introducing XTTS, which outperformed other systems with a score of 73.65, did not disrupt the relative ranking of the remaining systems, which remained consistent with the original MUSHRA test. Thus, there does not seem to be significant cognitive overload, as we still observe consistent results. Note that we study cognitive load using $n=7$ systems, but it remains to be seen how large $n$ can be before cognitive overload starts impacting the scores. For detailed scores, please refer to Table 5 in appendix.
@@ 5 Rethinking MUSHRA
We understand that devising a scoring formula involves some subjectivity. To address this, we encouraged raters to review their evaluations and adjust their fine-grained scores if they feel the overall scores from the formula did not accurately reflect the differences they perceived between system pairs. This way, the final ratings better represented the raters’ true opinions about each system’s quality and reduce any shortcomings that could have stemmed from the formula. More interestingly, we notice that the scores derived from the MUSHRA-DG test preserve the rankings obtained from the gold-standard Comparative Mean Opinion Score (CMOS) tests (Table 3), thus reinforcing the validity of our evaluations. Additionally, the variance in MUSHRA scores calculated using the formula is significantly lower, indicating reduced ambiguity in ratings and providing a clearer distinction between different systems’ performances.
@@ 6.1 Evaluations using MUSHRA-NMR
FS2 61.99 23.86 0.46 73.02 12.05 0.63 85.76 8.38 0.42
ST2 68.09 22.01 0.43 73.03 11.87 0.62 89.55 8.29 0.35
VITS 68.75 21.04 0.41 75.08 10.93 0.57 89.33 7.11 0.42
Anchor-X 71.83 19.97 0.39 77.67 14.60 0.76 89.73 6.96 0.38
Hindi Reference 76.39 18.08 0.35 91.45 10.30 0.54 89.45 7.47 0.36
Anchor-Y 21.94 16.74 0.38 45.63 10.14 0.35 56.01 7.46 0.34
FS2 66.77 19.12 0.35 82.32 6.93 0.39 78.01 6.79 0.32
VITS 68.52 18.28 0.36 82.32 5.60 0.28 78.33 6.39 0.32
ST2 76.64 17.68 0.33 88.50 7.72 0.51 88.77 6.33 0.38
Tamil Reference 78.69 17.26 0.34 93.18 7.34 0.37 95.47 10.62 0.54

### v2

@@ Abstract
Despite rapid advancements in TTS models, a consistent and robust human evaluation framework is still lacking. For example, MOS tests fail to differentiate between similar models, and CMOS’s pairwise comparisons are time-intensive. The MUSHRA test is a promising alternative for evaluating multiple TTS systems simultaneously, but in this work we show that its reliance on matching human reference speech unduly penalises the scores of modern TTS systems that can exceed human speech quality. More specifically, we conduct a comprehensive assessment of the MUSHRA test, focusing on its sensitivity to factors such as rater variability, listener fatigue, and reference bias. Based on our extensive evaluation involving 492 human listeners across Hindi and Tamil we identify two primary shortcomings: (i) reference-matching bias, where raters are unduly influenced by the human reference, and (ii) judgement ambiguity, arising from a lack of clear fine-grained guidelines. To address these issues, we propose two refined variants of the MUSHRA test. The first variant enables fairer ratings for synthesized samples that surpass human reference quality. The second variant reduces ambiguity, as indicated by the relatively lower variance across raters. By combining these approaches, we achieve both more reliable and more fine-grained assessments. We also release Mango, a massive dataset of 246,000 human ratings, the first-of-its-kind collection for Indian languages, aiding in analyzing human preferences and developing automatic metrics for evaluating TTS systems.
@@ 1 Introduction
The above problem is well studied for the Mean Opinion Scores (MOS) test (Wester et al., 2015; Finkelstein et al., 2023; Kirkland et al., 2023; Le Maguer et al., 2024) which has received much constructive criticism over the past few years. Specifically, in a MOS test, listeners assess each system independently, which can result in an inability to accurately capture the subtle relative differences between similar systems. This poses a significant challenge in modern TTS evaluation where systems that perform equally well need to be compared against each other. To address these issues some of the recent works rely on CMOS tests (Loizou, 2011). However, this test is costly and time-consuming as it involves ($N\choose 2$ comparisons between all pairs of $N$ systems.
The MUSHRA test has been gaining popularity in addressing these issues. This test scales better by enabling a parallel comparison of the $N$ systems, and addresses the limitations of MOS tests that only allow isolated evaluation. However, we show that even the MUSHRA test is not devoid of issues. To begin with, we note that the MUSHRA test was conventionally designed to assess intermediate-quality audio systems (ITU-R, 2015). However, state-of-the-art TTS systems (Ju et al., 2024) are not of intermediate quality and instead generate audios having quality on par or even better than human recordings. To align with these modern developments, several works adopt variants of MUSHRA (Merritt et al., 2022; Li et al., 2023; Shen et al., 2024), which differ in implementation but the validity of these modified tests is unknown.
@@
With the goal of seeking answers to the above questions, we collected 246,000 human ratings by conducting the MUSHRA test involving 3 systems across two languages, viz., Tamil and Hindi. Our in-depth analysis based on these ratings, reveals two primary shortcomings: (i) reference-matching bias and (ii) judgement ambiguity. To mitigate these issues, we propose two refined variants of the MUSHRA test. The first variant does not explicitly identify the human reference to the rater. Doing so, prevents unfair penalties for well-synthesized samples that differ from the human reference, such as those with natural prosody that do not match the reference’s prosody. In the second variant, raters are provided scoresheets to systematically calculate MUSHRA scores, by explicitly marking pronunciation mistakes, unnatural pauses, digital artifacts, word skips, liveliness, voice quality, rhythm, etc. Using the scores for these fine-grained criteria, they arrive at the final MUSHRA score. Our studies show that both these variants lead to a more reliable evaluation with the second variant also allowing for fine-grained fault isolation during evaluation. While MUSHRA-DG does require additional time to complete the tests, we believe that, given the limitations of the current MUSHRA setup, a slightly more time-intensive solution is justified to ensure the integrity and reliability of the evaluation process. We then show that a combination of these two approaches that leverages their individual strengths ensures both consistency and granularity. It allows modern TTS systems to be evaluated without being unfairly penalized for surpassing the reference in naturalness or prosody. The detailed scoring for pronunciation, prosody, and other factors provides actionable insights, and helps practitioners understand precisely where their systems excel and where improvements are needed. This combination creates a more balanced and sensitive evaluation framework, offering a clearer and more reliable assessment of TTS system performance.
@@
    Large-scale empirical validation of proposed variants resulting in Mango, a dataset of 246,000 ratings from 492 listeners across Hindi and Tamil, examining three TTS systems.
@@ 3 MANGO: A Corpus of Human Ratings for Speech
We introduce a new dataset, Mango: MUSHRA Assessment corpus using Native listeners and Guidelines to understand human Opinions at scale. It is a first-of-its-kind collection for any Indian language, comprising 246,000 human ratings of TTS systems and ground-truth human speech in both Hindi and Tamil, making it an expensive endeavour but one that we hope will contribute meaningfully to research in speech evaluation. Given the shortcomings of Mean Opinion Score (MOS) and Comparative Mean Opinion Score (CMOS) tests, our goal is to critically examine a promising alternative—the MUSHRA test—by conducting a large-scale evaluation involving multiple raters, systems, and languages. To do so, we adopt the standard MUSHRA test (ITU-R, 2015). Raters evaluate multiple stimuli on each page, including an explicit (mentioned) reference that serves as a benchmark for high-quality speech, along with an anchor and implicit (hidden) reference to calibrate judgments. Each stimulus is rated on a continuous scale from 0 to 100, which is also discretized: 100-80 (Excellent), 80-60 (Good), 60-40 (Fair), 40-20 (Poor), and 20-0 (Bad). We describe our evaluation setup below, and provide the detailed instructions provided to participants in Appendix A.1.
@@ 3.2 Synthesizing Speech Samples for Annotation
To generate samples for TTS evaluation, we train TTS systems on the Hindi and Tamil subsets of the IndicTTS database (Baby et al., 2016). Each language consists of recordings from a female and male speaker (Hindi: 20.17 hours; Tamil: 20.59 hours). We train FastSpeech2 (FS2) (Ren et al., 2021) with HiFiGAN v1 (Kong et al., 2020) and VITS (Kim et al., 2021) from scratch on the train-test splits using hyper-parameters suggested in a recent study (Kumar et al., 2023). We finetune StyleTTS2 (ST2) (Li et al., 2023) from the LibriTTS checkpoint.
@@ 3.3 Annotation Process and Dataset Statistics
With the above process we collected 246,000 human ratings for TTS systems. Table 1 shows the demographic distribution, the number of participants and the overall number of ratings across all MUSHRA tests and our proposed variants (MUSHRA-NMR, MUSHRA-DG, MUSHRA-DG-NMR) which are described in Section 5.
@@ 4.1 Is MUSHRA A Reliable Test?
In Table 3, we present the results of the MUSHRA test among 3 systems and find VITS and ST2 score highest in Hindi and Tamil respectively. Surprisingly, all systems attain scores in the “Good” bin with MUSHRA scores between 60 and 80, while the reference surpasses all systems with scores in the “Excellent" bin. Given that state-of-the-art TTS systems are able to reach quality on par with references, one would expect a much smaller gap between the reference and systems. To confirm this, we conduct the more reliable but expensive CMOS test with 15 listeners in each language. In this test, we ask the rater to compare a given system, such as VITS, with a reference audio sample. The rater evaluates both the reference and the output from a system being tested without prior knowledge of which audio sample corresponds to which system, ensuring an unbiased comparison. The raters assign a single score ranging from -3 to +3 in increments of 0.5. A score of -3 indicates that System A is much worse than System B . A score of +3 indicates that System A is much better than System B. A score of 0 means that both systems are equal in quality. As seen from the scores in Table 3, CMOS indicates that the outputs synthesised by VITS and ST2 are very close in quality to the reference in Hindi and Tamil respectively, while MUSHRA scores do not reflect this at all. We hypothesize that listeners in the MUSHRA test are subject to various biases, one of which we term the reference-matching bias. This bias may lead to situations where systems that perform comparably to or better than the reference are rated less favorably, as listeners tend to focus on aligning their ratings with the reference outputs while evaluating the systems. While this may have been acceptable when TTS systems lagged behind human speech quality, it is undesirable in the current scenario where modern TTS systems often exceed the reference in aspects like naturalness and prosody Li et al. (2023); Shen et al. (2024). This suggests that the MUSHRA test, in its conventional form, may no longer be sufficient for evaluating state-of-the-art TTS systems. Instead, alternative methodologies, such as the variants we propose in Section 5, may help ensure more fair and accurate assessments.
@@ 4.2 How reliable is the mean statistic in MUSHRA scores?
To delve deeper into judgment ambiguity, we examine variations between two systems. We consider an utterance where the mean scores for the samples generated by VITS and ST2 are nearly identical, but the variance across raters for each system is high. This high variance indicates significant ambiguity. Upon listening to many such utterances and speaking to many raters, we hypothesize that the ambiguity likely stems from different raters focusing on different aspects of the generated samples. For instance, some raters may prioritize prosody, others voice quality, and yet others the presence of digital artifacts. We hypothesize that asking raters to highlight these subtle differences across multiple dimensions while assigning a single score can lead to ambiguity in determining how much to penalize or reward a system’s output. Hence, clear guidelines which take into account a fine-grained evaluation across different aspects would help (as proposed later in Section 5).
@@ 4.4 What is the impact of rejecting raters per standard MUSHRA protocol?
Traditionally, MUSHRA employs a rater rejection criterion, wherein raters scoring the hidden reference (HR) below a threshold ($\lambda$) more than 15% of the time are rejected. This rejection rule stems from the inherent assumption that the HR is the gold standard, which is not true in modern TTS settings where TTS systems (Li et al., 2023; Shen et al., 2024) that is either on par with or surpasses the reference. In such cases, a rater consistently scoring the HR lower might not necessarily reflect unreliability, but rather a nuanced perception of the reference’s limitations compared to the evaluated systems. This is reinforced by observations from Table 3 where raters clearly prefer ST2 over reference for Tamil with a mean CMOS score of 0.24. This observation is further supported by the MUSHRA scores shown in Figure 3. The table shows that while rejecting raters based on the conventional MUSHRA criterion does not affect system rankings, it does notably shift scores. Specifically, system scores decrease with increasing $\lambda$, while reference scores increase. This trend hints at the reference-matching bias wherein raters who give the HR high scores might be unconsciously matching system samples to the mentioned reference, rather than rating based on their absolute perception of quality.
@@ 4.6 Does adding more systems affect scores?
To better understand cognitive overload in the MUSHRA test, we scale up the number of systems to be rated by introducing one new competitive system - XTTSv2, and repeating an existing system - VITS (VITS-R) in the original Hindi MUSHRA test. We finetune XTTSv2 (CoquiAI, 2023) starting from the multilingual checkpoint with the hyper-parameters from their original implementations on the same splits described in Section 3.2. We call this test - MUSHRA-Extended. The results show that raters were highly consistent, with VITS and VITS-R receiving nearly identical scores (68.99 and 68.47, respectively), despite the randomized order. Introducing XTTS, which outperformed other systems with a score of 73.65, did not disrupt the relative ranking of the remaining systems, which remained consistent with the original MUSHRA test. Thus, there does not seem to be significant cognitive overload, as we still observe consistent results. Note that we study cognitive load using $n=7$ systems, but it remains to be seen how large $n$ can be before cognitive overload starts impacting the scores. For detailed scores, please refer to Table 4.
Table 4: MUSHRA-Extended scores with 95% CI for Hindi.
System $\mu$ $\sigma$
CI
FS2 63.12 21.30 0.93
ST2 65.15 21.76 0.95
VITS-R 68.47 19.70 0.86
VITS 68.99 19.67 0.86
ANC 73.62 19.56 0.79
XTTS 73.65 18.52 0.86
REF 76.39 18.05 0.81
@@ 5 Rethinking MUSHRA
We understand that devising a scoring formula involves some subjectivity. To address this, we encouraged raters to review their evaluations and adjust their fine-grained scores if they felt the overall scores from the formula did not accurately reflect the differences they perceived between system pairs. Notably, we tracked these revisions and found that they occurred in only 1.3% of cases. This way, the final ratings better represented the raters’ true opinions about each system’s quality while also reducing any shortcomings that could have stemmed from the formula. More interestingly, we notice that the scores derived from the MUSHRA-DG test preserve the rankings obtained from the gold-standard Comparative Mean Opinion Score (CMOS) tests (Table 3), thus reinforcing the validity of our evaluations. Additionally, the variance in MUSHRA scores calculated using the formula is significantly lower, indicating reduced ambiguity in ratings and providing a clearer distinction between different systems’ performances.
@@ 6.1 Evaluations using MUSHRA-NMR
FS2 61.99 23.86 0.46 72.73 11.65 0.51 81.51 11.50 0.50
ST2 68.09 22.01 0.43 73.41 12.03 0.53 84.97 12.22 0.54
VITS 68.75 21.04 0.41 75.62 10.97 0.48 85.68 10.29 0.45
Anchor-X 71.83 19.97 0.39 80.67 13.57 0.59 88.45 7.38 0.32
Hindi Reference 76.39 18.08 0.35 90.87 9.34 0.41 88.63 7.39 0.32
Anchor-Y 21.94 16.74 0.38 45.00 11.47 0.35 54.66 17.91 0.46
FS2 66.77 19.12 0.35 82.36 8.06 0.32 82.87 10.48 0.35
VITS 68.52 18.28 0.36 81.96 7.41 0.35 83.86 10.12 0.44
ST2 76.64 17.68 0.33 88.82 7.88 0.50 91.10 8.05 0.78
Tamil Reference 78.69 17.26 0.34 94.61 6.98 0.31 95.99 6.86 0.30